OpenStudy (anonymous):

find a formula for the inverse of the function f(x)=(x^3)+1

OpenStudy (campbell_st):

for the inverse, just swap x and y then make y the subject. so after the swap you have \[x = y^3 + 1\] make y the subject.

OpenStudy (anonymous):

what do I do after \[x-1=y ^{3}\]

OpenStudy (campbell_st):

whats the opposite of cubing a number...?

OpenStudy (anonymous):

cubed root?

OpenStudy (campbell_st):

yes thats it... so take the cube root of both sides of the equation

OpenStudy (anonymous):

so would the answer be \[\sqrt[3]{x-1}=y\]

OpenStudy (campbell_st):

yes, great job

OpenStudy (anonymous):

ok thanks!